OpenStudy (anonymous):

Find the Product (8x-1/2)(8x+1/2) (Eight x minus one half)(eight x plus one half)

OpenStudy (oaktree):

This is a difference of two squares factorization.

OpenStudy (oaktree):

You need to square the first and last terms only.

OpenStudy (oaktree):

(8x)^2=64x^2 (1/2)^2 = 1/4 So, the multiplied binomial is 64x^2-1/4

OpenStudy (anonymous):

Thank you sooooo much!:)